what are the x and y intercepts of
8x+6y=13

Answer by Alan3354(69443)

what are the x and y intercepts of
8x+6y=13
-------------
The y-axis is crossed at x = 0
8*0 + 6y = 13
y = 13/6
--> (0,13.6)
-------------
The x-axis is crossed at y = 0
Do it the same way.